the water can be made to boil at 0 deg C if pressure of surrounding is______mm of Hg


a)76


b)5 cm


c)0.1


d)4.6


PLS TEL BE IF THERE IS ANY FORMULA FOR THIS>> IF ANY <<

Grade:9

1 Answers

Chetan Mandayam Nayakar
12 years ago

boiling point(in fahrenheit) = 49.161 * Ln (Pressure in inches Hg) + 44.932


Ln (Pressure in inches Hg)=32-44.932= -12.932,
(Pressure in inches Hg)=e-12.932 =2.42*10-6=2.42*2.54*10-5 =6.1468*10-5 mm Hg